# A Computational Theory for Efficient Model Evaluation with Causal Guarantees

### v3

@@ 5 Assessing assumptions
The IRIS assumption can be satisfied in data collection while the other assumptions may not be satisfied in real world. In order to assess the IIDE and unbiased assumptions, statistical analyses can be employed to verify necessary conditions under those assumptions, providing credibility to the theorems derived from this assumption for someone. It is important to note that these assessments do not conclusively prove the unbiasedness assumption. Instead, they serve as methods to potentially refute the assumption under specific conditions, rather than confirming its validity.
@@ 5.2 Assessing Unbiased assumption
Our bounds requires unbiased evaluation models while sometime the evaluation model may not be unbiased. To assess the unbiasedness of the evaluation model, we employ the following statistical procedures: 1) Global Unbiasedness Assessment: We conduct a one-sample t-test on the entire dataset to determine if the model’s errors have a mean of zero, indicating unbiasedness across the universal set; 2) Subset Unbiasedness Assessment: We randomly select 30 subsets from the dataset and perform one-sample t-tests on each to evaluate whether the mean errors within these subsets deviate significantly from zero. This step examines the consistency of the model’s unbiasedness across different data partitions.
@@ 6 Experiment
We performed experiments for evaluation model without condition and evaluation model with condition respectively in 12 scenes of model evaluation, including randomized experiment prediction, scientific simulation, advertising, insurance, and quantum trade. The detailed evaluation scenes (including experiment setting), assumption assessing, hyperparameter tuning, and evaluation results are listed in the appendix B, C, D, E.
@@ 6.1 Evaluation model
To demonstrate the generalizability of our evaluation model in subject space, we consider 11 distinct scenes. In these scenes, computational evaluation can be helpful to reduce experimental cost, shorten simulation time, or enhance the efficiency of advertising and insurance sales.
Although subjects are independently, randomly, and identically sampled from the subject space, they cannot be randomly deployed in real-world settings for direct measurement of benefit-related metrics (e.g., mortality, rehabilitation time). Instead, our evaluation system defines the evaluation metric as the post-experiment metric on unseen test data with randomized treatment. This metric captures the causal effect from subject to post-experiment outcome, as established in Theorem 4. Empirical studies (Ai et al. (2022); Zhou et al. (2023, 2024a); He et al. (2024); Gentzel et al. (2021)) have demonstrated a strong correlation between the decision model’s predictive performance on randomized experimental data and its performance following randomized deployment. Accordingly, we employ these predictive metrics to approximate randomized post-deployment performance, which can then be integrated into the decision via an appropriate utility function.
@@
Across the 11 scenes, we compare our evaluation model against six existing validation methods that do not incorporate prior information from data. The normalized negative empirical RMSE is illustrated in Figure 3. Our best evaluation model achieves error reductions of 94.7–99.0% in RMSE, 81.7–95.5% in $R^{2}$, 24.1–77.0% in ROC-AUC, and 39.2–89.5% in ACC relative to the holdout-100 validation method. This performance improvement is attributable to the effective modeling of the relationships among the subject, its proxy metrics, and the evaluation metrics. Notably, models that capture spurious correlations (MLP and SVR without hyperparameter tuning) fail to reduce evaluation error.
@@ 6.4 Evaluation cost
The acceleration ratio of time achieved by our computational method ranges from 1,000 to 10,000,000 per evaluation subject in the 12 scenes as shown in appendix. Although experimental and simulation-based evaluations can be accelerated via parallel processing (excluding additional evaluation costs), the same parallelization strategies can be readily applied to computational evaluation approach. The detailed comparison of evaluation time can be seen in the appendix F.
@@ 7 Limitation and conclusion
Limitation. Despite the theoretical contributions and empirical results, the proposed approach is subject to several limitations. First, our theoretical guarantees rely on the IIDE, IRIS, and unbiasness assumptions, which may not hold in all real-world scenes although statistical tests provide some support. Second, while our method has been validated across multiple application domains, its scalability and generalization in extremely high-dimensional settings require further investigation. Third, lower error and tighter bound are required to apply evaluation models into real applications with small sample.
@@
Conclusion. In this work, we propose a novel computational framework for the evaluation of decision models that rigorously derives upper bounds on generalized evaluation error and causal effect error, and employs a meta-learning strategy to address heterogeneity in evaluation subject spaces. Our extensive experimental results, conducted over 12 diverse scenes—including individual medicine, scientific simulation, social experiments, business activity, and quantum trade—demonstrate significant error reductions (ranging from 24.1% to 99.0%) and substantial acceleration (up to $10^{7}$) compared to traditional evaluation methods. These findings highlight the potential of our technical route to drastically reduce evaluation costs while enabling rapid model iteration. Future work will focus on scaling the framework to larger domains and enhancing the validation of key assumptions to further improve robustness and applicability. The paradigm of computational evaluation can be seen in the appendix H.
Appendix A Proofs
@@
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled from distribution $Pr(S)$, $m$ is generate by a unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is root mean square error, $\forall f_{1}$, $\forall f_{2}$ where $f_{1}$ and $f_{2}$ are unbiased estimation of $f$ given $c$, we have Efficiency: if $E_{gen}(f_{1})<E_{gen}(f_{2})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
@@
$\forall k=1,2$, denote $\epsilon_{k}(c,s,e)=f_{k}(c,s)-f(c,e,s)$, according to conditional unbiasness, we have $E(\epsilon_{k}(c,s))=0$. The error is denoted as
$\Delta f_{k}(s)=f_{k}(c,s_{a})-f_{k}(c,s_{b})-f(c,s_{a},e)-f(c,s_{b},e)=\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e)$
@@
$(\Delta f_{k})^{2}=(\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e))^{2}$
@@
$E((\Delta f_{k})^{2})=E(\epsilon_{k}(c,s_{a},e)^{2})+E(\epsilon_{k}(c,s_{b},e)^{2})-2E(\epsilon_{k}(c,s_{a},e)\epsilon_{k}(c,s_{b},e))$
. According to the condition unbiasness and IRIS assumption, the third term
@@
, then because $\forall s$, $E_{gen}(f_{1})<E_{gen}(f_{2})$, that is $E(\epsilon_{1}(c,s,e)^{2})<E(\epsilon_{2}(c,s,e)^{2})$ for any $s$, then
$E((\Delta f_{1})^{2})=E(\epsilon_{1}(c,s_{a},e)^{2})+E(\epsilon_{1}(c,s_{b},e)^{2})<E(\epsilon_{2}(c,s_{a},e)^{2})+E(\epsilon_{2}(c,s_{b},e)^{2})=E((\Delta f_{2})^{2})$
, which can be write as $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$ when error function is mean square error. Following, the consistency can be proved spontaneously by the Squeeze Theorem in Calculus. ∎
@@
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, randomly, identically sampled error measurements where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error.
@@
. Due to IIDE and IRIS assumption, the sub population of errors where $s=a$ and $s=b$ is also independently sampled,
@@
. Due to the unbiased assumption and IIDE assumption for populations where $s=s_{a}$ and $s=s_{b}$,
@@
. Denote $\epsilon_{0}=2\max\{\nu_{a},\nu_{b}\}$ where $\nu_{a}=E_{emp}(a)+\sqrt{\frac{1}{2n_{a}}\ln(\frac{2}{\sigma}})$ and $\nu_{b}=E_{emp}(b)+\sqrt{\frac{1}{2n_{b}}\ln(\frac{2}{\sigma})}$, due to Bonferroni inequality,
@@
Causal bound with non-positivity. Given arbitrary conditional unbiased (CU) evaluation model $f$ given any $c$, then $\forall s_{a}\in S,\forall s_{b}\in S$,
@@
. Due to IIDE and IRIS assumption, the sub population of errors where $s=a$ and $s=b$ is also independently, randomly and identically sampled following the whole population,
@@
$P(E_{gen}(\Delta\hat{f})\leq 2(E_{emp}(\Delta\hat{f})+\sqrt{\frac{1}{2n}\ln(\frac{1}{\sigma})}))\geq 1-\sigma$
Appendix B Evaluation scenes
We introduce 12 scenes to test the performance of our (conditional) evaluation model. The evaluation subject and evaluation metric in those applied scenes is listed in the table 3.
The heterogeneous evaluation subject space is configured as a combination of a linear subject space (logistic regression or linear regression) and a non-linear subject space (MLP Classifier or MLP Regressor). The model’s input dimension is the number of features, and output dimension is the number of the targets. The MLP’s hidden layer size limit is set as 8.
The IRIS sampling of the subject is two steps. The first step is the subject type sampling with 0.5 probability to choose linear model and 0.5 probability to choose MLP model. The second step is to choose a specific model in the sub space. For both linear model and MLP, we use Normal distribution $N(0,1)$ to determine the parameter of the models. The outcome when we calculate Shapley value is set as the negative evaluation errors.
Scene Evaluation Subject (Model) Application Metric

### v4

@@ 5 Assessing assumptions
The IRIS assumption can be satisfied in data collection while the other assumptions may not be satisfied in real world. While the IIDE assumption could be trivially met by employing a random number generator as the outputs for IID outcome, this approach is devoid of practical utility, as it fails to reflect any substantive patterns or domain-specific knowledge required for realistic evaluation modeling. In order to assess the IIDE and unbiased assumptions, statistical analyses can be employed to verify necessary conditions under those assumptions, providing credibility to the theorems derived from this assumption for someone. It is important to note that these assessments do not conclusively prove the assumptions. Instead, they serve as methods to potentially refute the assumptions under specific conditions, rather than confirming its validity.
@@ 5.2 Assessing Unbiased assumption
Our effect error bounds requires unbiased evaluation models while sometime the evaluation model may not be unbiased. To assess the unbiasedness of the evaluation model, we employ the following statistical procedures: 1) Global Unbiasedness Assessment: We conduct a one-sample t-test on the entire dataset to determine if the model’s errors have a mean of zero, indicating unbiasedness across the universal set; 2) Subset Unbiasedness Assessment: We randomly select 30 subsets from the dataset and perform one-sample t-tests on each to evaluate whether the mean errors within these subsets deviate significantly from zero. This step examines the consistency of the model’s unbiasedness across different data partitions.
@@ 6 Experiment
We performed experiments for evaluation model without condition and evaluation model with condition respectively in 12 scenes of mini agent evaluation, including randomized experiment prediction, scientific simulation, advertising, insurance, and quantum trade. The detailed evaluation scenes (including experiment setting), assumption assessing, hyperparameter tuning, and evaluation results are listed in the appendix D, E, F, G.
@@ 6.1 Evaluation model
To demonstrate the generalizability of our evaluation model in agent space, we consider 11 distinct scenes. In these scenes, computational evaluation can be helpful to reduce experimental cost, shorten simulation time, or enhance the efficiency of advertising and insurance sales.
Although subjects are independently, randomly, and identically sampled from the subject space, they cannot be randomly deployed in real-world settings for direct measurement of benefit-related metrics (such as mortality, rehabilitation time) due to experimental limitation. Instead, our evaluation system defines the evaluation metric as the post-experiment metric on unseen test data with randomized treatment. The metrics capture the causal effect from subject to post-experiment outcome, as established in Theorem 4. Empirical studies (Ai et al. (2022), Zhou et al. (2023), Zhou et al. (2024a), Cheng et al. (2022), Gentzel et al. (2021)) have also demonstrated the correlation between the agents’ predictive performance on randomized experimental data and its performance following randomized deployment. Accordingly, we employ these predictive metrics to approximate randomized post-deployment performance, which can then be integrated into the AI decision via an appropriate utility function.
@@
Across the 11 scenes, we compare our evaluation model against 6 existing validation methods that do not incorporate prior information from data. The normalized negative empirical RMSE is illustrated in Figure 3. Our best evaluation model (Het-Linear) achieves error reductions of 94.7% to 99.0% in RMSE, 81.7% to 95.5% in $R^{2}$, 24.1% to 77.0% in ROC-AUC, and 39.2% to 89.5% in ACC relative to the Holdout-100 validation method. This performance improvement is attributable to the effective modeling of the relationships among the subject, its proxy metrics, and the evaluation metrics. Notably, models that capture spurious correlations (MLP and SVR without hyperparameter tuning) fail to reduce evaluation error.
@@ 6.4 Evaluation cost
The acceleration ratio of time achieved by our computational method ranges from 1,000 to 10,000,000 per evaluation subject in the 12 scenes as shown in appendix H. Although experimental and simulation-based evaluations can be accelerated via parallel processing (excluding additional evaluation costs), the same parallelization strategies can be readily applied to computational evaluation approach.
@@ 7 Limitation and conclusion
Limitation. First, our theoretical guarantees rely on three assumptions, which may not hold in all real-world scenes. Second, while our method has been validated across multiple domains, its generalization in extremely high-dimensional settings (such as LLM) require further investigation. Third, lower error and tighter bound are required to apply evaluation models into real applications with small sample. Four, scalability of evaluation models across scenes is another problem in future study. Five, the post-experiment metric (such as mortality, RoI) of randomized deployed mini-agents with high utility should be collected to increase the value of computational evaluation.
@@
Conclusion. In this work, we propose a novel computational framework for the evaluation of mini-agents that rigorously derives upper bounds on generalized evaluation error and effect error, and employs a meta-learning strategy to address heterogeneity in agent space. Our extensive experimental results, conducted over 12 diverse scenes—including individual medicine, scientific simulation, social experiments, business activity, and quantum trade—demonstrate significant error reductions (ranging from 24.1% to 99.0%) and substantial acceleration (up to $10^{7}$) compared to traditional evaluation methods. These findings highlight the potential of our technical route to drastically reduce evaluation costs for rapid model iteration. Future work will focus on benefit-related metrics and enhancing the validation of key assumptions for applicability. The upper bound table and paradigm of computational evaluation were given in the appendix I and J.
Appendix A Investigation of individual medical AI in randomized controlled trial
From January 2022 and January 2023, there are total 380 completed phase III interventional trials (adult) with results and study documents on ClinicalTrials.gov. The statistical estimates can be categorized into six classes: geometric mean, mean, square mean, median, rank, and percentage, as shown in the figure 5. It is important to note that these estimates can not reflect the probability of a group or an individual benefiting from the treatment. For example, for treatment effect with Gaussian distribution, the mean difference can still be significant even the treatment is harmful for $49\%$ and benefited for $51\%$ if $n$ is large enough. The randomized controlled trails can not be applied to development of individual treatment AI directly.
Figure 5: Estimates percentage of completed phase III interventional trials (adult) with results and study documents on https://clinicaltrials.gov.
Another query was conducted on ClinicalTrials.gov using the keyword “Artificial Intelligence” under the Condition or Disease field. Filters were applied to select only completed, interventional studies, with posted results, and available statistical analysis plans, with no restriction on completion date at January of 2025.
Only a total of 13 studies were retrieved. After manually excluding single-arm trials (i.e., studies without control groups or randomized design), 7 randomized clinical trials (RCTs) (Nayak et al. [2023], Abramoff et al. [2023], Lv et al. [2023], MacNeill et al. [2024], Eng et al. [2021], Piette et al. [2022], Mohr et al. [2019]) remained that investigated the clinical efficacy of AI-based healthcare interventions.
Among these, only 3 trials reported statistically significant primary outcomes, including:
  • An AI-powered blood glucose reminder system (Nayak et al. [2023]),
  • The Wysa AI chatbot for mental health support (MacNeill et al. [2024]),
  • An AI system for bone age prediction (Eng et al. [2021]).
These results highlight the challenge of translating AI systems into demonstrably effective clinical interventions, emphasizing the need for rigorous evaluation, robust study design, and careful integration into complex healthcare workflows.
Appendix B Investigation of agent evaluation
B.1 Existed agent evaluation works
The theoretical foundations of agent evaluation, particularly with causal guarantees, remain underdeveloped. To date, we are not aware of any systematic theory that rigorously supports causal inference for agent performance evaluation. While several benchmarking efforts exist, they primarily focus on performance measurement rather than causal understanding. MLAgentBench (Huang et al. [2024]) uses 13 tasks (including Canonical Tasks, Classic Kaggle, Kaggle Challenges, Recent Research, and Code Improvement) to evaluate the performance of large language model. SUPER Benchmark (Bogin et al. [2024]) is used to evaluate the agents’ ability to reproduce results from research repositories. AI Clinician (Komorowski et al. [2018]) and AI Hospital (Fan et al. [2025]) use synthetic data to test the performance of clinical agents. Jia et al. [2024] explores the LLMs’ output decisions’ alignment with human norms and ethical expectations. However, a common limitation across these works is the lack of control over unobserved confounders, such as participants’ pre-experiment training or researchers’ design biases. These hidden common causes may simultaneously influence both the deployed agent and its measured performance in real-world settings. As a result, these benchmarks cannot identify the (conditional) causal effect of an agent on real-world outcomes—such as clinical mortality—limiting their utility for causal evaluation.
B.2 Causal mini agent evaluation and potential outcome
Despite targeting different problem settings, the causal computational evaluation framework and the potential outcomes framework (Imbens and Rubin [2015]) for agent evaluation share one key connection and exhibit two fundamental differences. The connection lies in the fact that the unconfoundedness (ignorability/exchangeability) assumption in potential outcomes can be directly implied by the IRIS assumption. Furthermore, IRIS can be satisfied in practice by deploying mini agents in a randomized manner. The first major difference is that the Stable Unit Treatment Value Assumption (no interference and consistency)—a common requirement in potential outcome formulations—is not necessary under our framework, as long as the IIDE (Independent and Identically Distributed Evaluations) assumption holds. The second key difference is that the positivity assumption, which typically requires a non-zero probability of receiving each treatment, is also not required. Nevertheless, the causal effect of mini agents on evaluation metrics can still be bounded in the limit, as formally established in Theorem 4. These distinctions suggest that the proposed framework can be viewed as a generalized and more flexible extension of the potential outcomes framework, tailored specifically for the causal evaluation of mini agent behaviors.
B.3 Existed causal inference benchmark and packages
Although a number of benchmark datasets and open-source packages have been developed to address general causal inference problems—such as CausalML (Zhao and Liu [2023]), EconML (Battocchi et al. [2019]), DoWhy (Blöbaum et al. [2024], Sharma and Kiciman [2020]), CauseBox (paras [2612]), CausalNex (quantumblacklabs ), Causal Curve (Kobrosly [2020]), CausalDiscovery (Kalainathan et al. [2020]), pcalg (Kalisch et al. [2012]), bnlearn (Scutari [2010]), TETRAD (Ramsey et al. ), CausalityBenchmark (Shimoni et al. [2018]), JustCause (inovex ), CausalEffect (Tikka and Karvanen [2017]), Ananke (Lee et al. [2023]), Dagitty (Textor et al. [2011]), YLearn (Bochen Lyu [2022]), CausalImpact (Google ), causal-learn (Zheng et al. [2024]), gCastle (Zhang et al. [2021]), dosearch (Tikka et al. [2021]), CEE (Li ), awesome-causality-identification (YAN [2025]) —these tools are primarily designed for causal discovery or treatment effect estimation. Despite extensive investigation, we did not identify any existing algorithms or frameworks that directly address the problem of causal evaluation of agents, particularly in the context of evaluating agent-driven decision-making processes.
Appendix C Proofs
@@
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled (IRIS) from distribution $Pr(S)$, $m$ is generate by an unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is mean square error, $\forall\hat{f_{1}}$, $\forall\hat{f_{2}}$ where $\hat{f_{1}}$ and $\hat{f_{2}}$ are unbiased estimation of $f$ given $c$ (CU), we have Efficiency: if $E_{gen}(\hat{f_{1}})<E_{gen}(\hat{f_{2}})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(\hat{f_{1}}(c,s_{a})-\hat{f_{1}}(c,s_{b}))<E_{gen}(\hat{f_{2}}(c,s_{a})-\hat{f_{2}}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
@@
$\forall k=1,2$, denote $\epsilon_{k}(c,s,e)=\hat{f_{k}}(c,s)-f(c,e,s)$, according to conditional unbiasness, we have $E(\epsilon_{k}(c,s))=0$. The error is denoted as
$\Delta\hat{f_{k}}(s)=\hat{f_{k}}(c,s_{a})-\hat{f_{k}}(c,s_{b})-(f(c,s_{a},e)-f(c,s_{b},e))=\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e)$
@@
$(\Delta\hat{f_{k}})^{2}=(\epsilon_{k}(c,s_{a},e)-\epsilon_{k}(c,s_{b},e))^{2}$
@@
$E((\Delta\hat{f_{k}})^{2})=E(\epsilon_{k}(c,s_{a},e)^{2})+E(\epsilon_{k}(c,s_{b},e)^{2})-2E(\epsilon_{k}(c,s_{a},e)\epsilon_{k}(c,s_{b},e))$
. According to the IRIS assumption and condition unbiasness assumption, the third term
@@
, then because $\forall s$, $E_{gen}(\hat{f_{1}})<E_{gen}(\hat{f_{2}})$, that is $E(\epsilon_{1}(c,s,e)^{2})<E(\epsilon_{2}(c,s,e)^{2})$ for any $s$, then
$E((\Delta\hat{f_{1}})^{2})=E(\epsilon_{1}(c,s_{a},e)^{2})+E(\epsilon_{1}(c,s_{b},e)^{2})<E(\epsilon_{2}(c,s_{a},e)^{2})+E(\epsilon_{2}(c,s_{b},e)^{2})=E((\Delta\hat{f_{2}})^{2})$
, which can be write as $E_{gen}(\hat{f_{1}}(c,s_{a})-\hat{f_{1}}(c,s_{b}))<E_{gen}(\hat{f_{2}}(c,s_{a})-\hat{f_{2}}(c,s_{b}))$ when error function is mean square error. Following, the consistency can be proved spontaneously by the Squeeze Theorem in Calculus. ∎
@@
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, randomly, identically sampled error measurements (IIDE) where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, evaluation error function $L$ is mean square error ranged from 0 to 1.
@@
. Due to IIDE assumption, the sub population of errors where $s=a$ and $s=b$ is also independently sampled,
@@
. Due to the unbiased assumption,
@@
. Denote $\epsilon_{0}=2\max\{\nu_{a},\nu_{b}\}$ where $\nu_{a}=E_{emp}(a)+\sqrt{\frac{1}{2n_{a}}\ln(\frac{2}{\sigma}})$ and $\nu_{b}=E_{emp}(b)+\sqrt{\frac{1}{2n_{b}}\ln(\frac{2}{\sigma})}$, due to Bonferroni inequality (Galambos [1977]),
@@
. Due to IIDE and IRIS assumption, the sub population of errors where $s=a$ and $s=b$ is also independently sampled (IIDE) identically following the whole population (IRIS),
@@
Appendix D Evaluation scenes
We introduce 12 scenes to test the performance of our (conditional) evaluation model: Alert (Wilson et al. [2021]), Withdraw (Wilson et al. [2023]), Grid-S (Schäfer et al. [2016]), Higgs-S (Baldi et al. [2014]), Insurance (Kumar [2023]), Climate-S (Lucas et al. [2013]), TERECO (Li et al. [2022]), Calculus (Kramer et al. [2023]), Ad (Gokagglers [2018]), kin8nm-S (Fischer [2022]), NuScale-S (Huu Tiep [2024]), and Quantum trade. The evaluation subject (agent) and evaluation metric in those applied scenes is listed in the table 4. All the data are collected from real experiments or simulation. As mentioned in the section 6, we use the prediction performance in randomized experiment to approximate the post-experiment performance of randomized deployed mini agents. We also listed the website and license of the mentioned asset in the following table 3.
Scenes Website License
Alert https://datadryad.org/dataset/doi:10.5061/dryad.4f4qrfj95 CC0 1.0
Withdraw https://datadryad.org/dataset/doi:10.5061/dryad.kh189327p CC0 1.0
Grid-S https://www.openml.org/search?type=data&status=active&id=44973 CC BY 4.0
Higgs https://www.openml.org/search?type=data&status=active&id=42769 Public
Insurance https://www.kaggle.com/datasets/arun0309/customer Unknown
Climate https://www.openml.org/search?type=data&status=active&id=40994 Public
TERECO https://datadryad.org/dataset/doi:10.5061/dryad.59zw3r27n CC0 1.0
Calculus https://datadryad.org/dataset/doi:10.5061/dryad.kkwh70s95 CC0 1.0
Ad https://www.kaggle.com/datasets/loveall/clicks-conversion-tracking Other
kin8nm https://www.openml.org/search?type=data&status=active&id=44980 Public
NuScale https://archive.ics.uci.edu/dataset/1091 CC BY 4.0
Table 3: Website and license of the mentioned asset.
The heterogeneous mini agent space is configured as a combination of a linear mini agent space (logistic regression or linear regression) and a non-linear mini agent space (MLP Classifier or MLP Regressor). The agent’s input dimension is the number of features, and output dimension is the number of the targets. The MLP’s hidden layer size limit is set as 8.
The IRIS sampling of the agent is two steps. The first step is the agent type sampling with 0.5 probability to choose linear model and 0.5 probability to choose MLP model. The second step is to choose a specific model in the sub space. For both linear agent and MLP agent, we use Normal distribution $N(0,1)$ to determine the parameter of the models. The outcome when we calculate Shapley value is set as the negative evaluation errors.
Scene Evaluation Subject (Agent) Application Metric